a bullet is fired with a speed of 1000 m/s in order to hit a target 100 m away. if g=10m/s2 the gun shoul be aimed..
a) directly towards target
b) 5 cm above the target
c) 10cm above the target
d) 15cm above the target

Grade:11

3 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Hi,

Time taken by bullet to cover 100 m horizontally = (100/1000) = 0.1 s
Distance covered by bullet vertically in that time = s = 0.5 * 10 * 0.12= 0.05 m
So, the bullet should be fired 5cm above the target.

thanks.
Lekshmi
6 years ago
Distance=100m
Speed=1000m/s
Time=100/1000
=0.1s
substitute t in eqn s=ut+0.5at^2
Since u is zero s=0.5at^2
=0.5*10*0.1^2
=0.05m
=5 cm
Since value is positive 
Gun must be aimed at 5cm above target
Option 2 is correct
Thank you.
 
Lavish
4 years ago
Speed of bullet= 1500m/s 
Distance of target =150m
So the time taking to hit the target t=distance/speed 
t=150/1500=0.1sec
Given that g=10m/sec^2
So the bullet attracted by gravity in 0.1 sec
Because of gravity--
so,In 1sec^2 the bullet went down = 10m.
and 0.1sec^2 the bullet went down=10×0.1sec^2
                                                                   =10×0.01
                                                                   =0.1m. OR 10cm. Down from the target.